Determine the rate of return for the following cash flow. Year نہ سا Net cash flow -80000 -9000 40000 70000

Solve using Excel please!
Answer #1

Net cash inflow = -80000-9000+40000+70000 = $21,000

Rate of Return = $21000 / 80000 = 26.25%